# Efficient Bayesian Computation Using Plug-and-Play Priors for Poisson Inverse Problems

### v1

@@ 4.2.2 Experimental results
A set of representative results with the leaves image from set3c is presented in Figure 5 (additional results are available in Appendix A). Figure 5(a) depicts a realization $y$ generated by the forward model Eq. (1) with the operator $A$ modelling a Gaussian blur operator of size $25\times 25$ with standard deviation $1.6$ pixels and a photon level set to $\alpha=20$. Figures 5(b) to 5(f) show the posterior mean $\hat{x}_{MMSE}$, as calculated for the different denoisers and different algorithms. We observe that Prox-DRUNet outperforms the other denoisers in terms of reconstruction quality for both Euclidean and non-Euclidean algorithms and see that both algorithms deliver highly similar estimates, which capture fine detail without noticeable artefacts. In comparison, the LMMO denoiser produces an estimate with excessive smoothing, while the GS-DRUNet exhibits stripe-like artefacts. These artefacts are also present when respecting the step size bounds recommended in [28] (see Section 3.1.2, not shown here), and increase gradually with greater step size. The B-DRUNet also exhibits artefacts, amplifying some structures in the top central part of the image.
@@
Lastly, Figure 6 shows the pixel-wise posterior standard deviation, as calculated with the LMMO, Prox-DRUNet, GS-DRUNet and B-DRUNet denoisers. For reference, next to each posterior standard deviation plot, we also report the residual obtained by comparing $\hat{x}_{MMSE}$ to the true image (note that these standard deviations represent the models’ marginal predictions for these residuals, at the pixel level). We observe that LMMO, Prox-DRUNet, GS-DRUNet produce uncertainty plots that are broadly in agreement with their residuals. For Prox-DRUNet and LMMO, uncertainty concentrates around edges and contours, as expected for a deconvolution problem, whereas the uncertainty estimates of GS-DRUNet are aligned with its reconstruction artefacts. Conversely, B-DRUNet produces uncertainty plots that highlight homogenenous regions, and which do not align well with its residual. We conclude that Prox-DRUNet is the most appropriate denoising architecture for Bayesian PnP inference in Poisson image deblurring problems and use this as the prior for the rest of our numerical experiments.
@@ 4.3.1 Experimental set up
We now compare different PnP Langevin strategies with Prox-DRUNet as prior. We study how different choices of step size affects convergence speed, reconstruction quality, and numerical stability, with the goal to explore the trade-off that yields the best results for each algorithm. We interpret loss of stability through artefacts that appear in the reconstructions. We compare MMSE reconstructions computed with different sampling algorithms for a range of step sizes, both quantitatively and qualitatively, to obtain step size recommendations for each algorithm. Additionally, we compare the convergence speed between different sampling algorithms for the recommended step size over a set of images by looking at mean cumulative quality metrics. How image quality and convergence speed interrelate for different algorithms can guide the choice of sampling algorithm according to given priorities with regards to metric and computational budget.
@@
For this experiment, we use images from the set3c data set, depicted in Figure 3, the blur kernel displayed Figure 2(b) and photon level $\alpha=20$. We choose the starfish image to illustrate qualitative results. With regards to the algorithms and their step sizes, for the RPnP-ULA (Algorithm 1) and PPnP-ULA (Algorithm 2), we test a range of step sizes $\delta=c\cdot\delta_{L}$, with $c\in[1,10^{3}]$. For RPnP-SKROCK (Algorithm 3), we test step sizes of the form $\delta=c\cdot\ell_{s}\delta_{L}$, with $c\in[0.1,10^{2}]$, where $\ell_{s}=(s-0.5)^{2}(2-4/3\eta)-1.5$ with $s=10$ and $\eta=0.05$; for more details see [43]. For the PnP-MLA (Algorithm 4), we test step sizes of the form $\delta=c\cdot\delta_{L}$, where $c\in[10,10^{3}]$. The Prox-DRUNet denoiser parameter $\epsilon$ is set to $\epsilon=(20/255)^{2}$ for RPnP-ULA, PPnP-ULA and RPnP-SKROCK, whereas we use $\epsilon=(25/255)^{2}$ for PnP-MLA (see Appendix B.2 for complementary results for PnP-MLA with $\epsilon=(20/255)^{2}$). In all cases, to improve numerical stability and reduce PnP artefacts, we enforce equivariance to horizontal and vertical flips as well as to rotations of multiples of 90 degrees by randomization.
@@ 4.3.2 Experimental results
In Figure 8.3, we observe that RPnP-SKROCK is stable as the step size increases and does not show artefacts in its reconstructions. For large step sizes, however, we see significant oversmoothing which is a sign of excessive bias, see e.g. Figure 8.3(d) and 8.3(e). In this case, PSNR and SSIM continue to improve, while LPIPS worsens, quantitatively confirming a loss of detail. The overall performance deteriorates for a step size for which $c>0.3\cdot 10^{2}$. Therefore, we recommend to choose a step size for which $c=0.02\cdot 10^{2}$ or $c=0.3\cdot 10^{2}$ depending on whether objective (PSNR, SSIM) or perceptive quality (LPIPS) is more important. In Figure 8.4, we observe that PnP-MLA recovers fine details without any oversmoothing effect for all the considered step sizes. Performance improves across all metrics as the step size increases, but only negligibly when $c>5\cdot 10^{2}$.
@@
Figure 10: Mean cumulative metrics for the set3c data set using selected time steps, with x-axis in log scale to show the tail behaviour more clearly.
Finally, in Figure 10 we compare quality metrics (PSNR and LPIPS) against neural function evaluations (NFE) averaged over set3c using recommended step sizes to assess the convergence speed. We observe that RPnP-SKROCK outperforms all other algorithms in terms of PSNR in performance and convergence speed. In addition, RPnP-ULA and PPnP-ULA show similar performance and convergence speed while they outperform RPnP-SKROCK and PnP-MLA in terms of LPIPS. This occurs due to RPnP-SKROCK and PnP-MLA introducing more bias by taking larger step sizes. PnP-MLA shows slower convergence in terms of PSNR but it performs similarly with the other algorithms in terms of LPIPS with sharp convergence in $\sim 10^{4}$ NFEs.
@@ 5 Conclusion
This paper presents two generalizations of the PnP-ULA framework of [28] for Poisson imaging problems: PnP Langevin methods derived from a reflected and regularized overdamped Langevin diffusion, and PnP mirror Langevin methods. We focus on two novel algorithms: RPnP-SKROCK, which is an accelerated PnP Langevin sampling method based on boundary reflections and an approximation of the Poisson likelihood function; and PnP-MLA, the first PnP mirror sampling algorithm in the literature, which exploits a change of geometry to simultaneously deal with the constraints and with the Poisson likelihood without approximations. Moreover, using Poisson image deblurring as testbed, we compared a wide range of image denoising architectures and algorithmic choices, and concluded that the Prox-DRUNet denoiser is a robust image prior for the considered imaging problems. Extensive numerical experiments show that the proposed sampling methods are particularly relevant in low-photon settings (small $\alpha$). In such cases, the proposed sampling algorithms outperform PnP optimization-based strategies in reconstruction accuracy, while also providing uncertainty quantification. Among these, PnP-MLA consistently achieves the highest perceptual quality (LPIPS), demonstrating superior detail recovery at low photon levels. This suggests strong potential for applications requiring faithful reconstruction of small structures in sparse photon-starved data. Alternatively, RPnP-SKROCK achieves remarkable MMSE performance (PSNR) with a comparatively low computational cost, requiring in the order of $5\cdot 10^{3}$ NFEs or less, while offering a similarly detailed uncertainty quantification. As a general rule, we would recommend PnP-MLA when $\alpha<20$ (for images with pixel values in the range $[0,1]$ and assuming $A$ is normalized), and otherwise RPnP-SKROCK.
@@
As mentioned previously, a main reason for constructing Markov kernels such as PnP-MLA and RPnP-SKROCK is to embed them within empirical Bayesian strategies that automatically calibrate unknown model parameters in semi-blind imaging problems (see, e.g., [64, 38]). This remains a main direction for future work. In addition, while in this paper we have focused on relatively conventional PnP denoiser architectures, we see great potential in using a PnP prior derived from a denoising diffusion model (DM) within a Langevin sampling scheme, like in [37]. Extending the DM-based PnP-ULA of [37] to Poisson imaging problems by using either PnP-MLA or RPnP-SKROCK is also an important perspective for future work. Lastly, a highly promising but more challenging direction for further research would be to explore accelerated PnP-MLA sampling, combining the strengths of PnP-MLA and RPnP-SKROCK.
@@
Figures 13(b) to 13(f) show the posterior mean $\hat{x}_{MMSE}$, as calculated for the different denoisers and different algorithms using $10^{6}$ iterations to stress-test the stability of the algorithms. Prox-DRUNet proved to be the most competitive in terms of stability and performance among all the considered denoisers. More specifically, we observe that PPnP-ULA with the LMMO denoiser produces an estimate of the mean with excessive smoothing, while with the GS-DRUNet exhibits checkerboard-like and stripe-like reconstruction artefacts on large parts of the starfish and the background. The B-DRUNet denoiser progressively develops artefacts after approximately $10^{5}$ iterations which deteriorate gravely over time, see Figure 13(f).

### v2

@@ 4.2.2 Experimental results
We performed a thorough evaluation of the denoisers on the set3c dataset. For clarity and detailed analysis, we present a representative set of results using the leaves image in the main text (Figure 5). To demonstrate the consistency of our findings, comprehensive results for the starfish image are provided in Appendix A. The results for the butterfly image were qualitatively similar and offered no new insights, and were thus omitted for brevity. Figure 5(a) depicts a realization $y$ generated by the forward model Eq. (1) with the operator $A$ modelling a Gaussian blur operator of size $25\times 25$ with standard deviation $1.6$ pixels and a photon level set to $\alpha=20$. Figures 5(b) to 5(f) show the posterior mean $\hat{x}_{MMSE}$, as calculated for the different denoisers and different algorithms. We observe that Prox-DRUNet outperforms the other denoisers in terms of reconstruction quality for both Euclidean and non-Euclidean algorithms and see that both algorithms deliver highly similar estimates, which capture fine detail without noticeable artefacts. In comparison, the LMMO denoiser produces an estimate with excessive smoothing, while the GS-DRUNet exhibits stripe-like artefacts. These artefacts are also present when respecting the step size bounds recommended in [35] (see Section 3.1.2, not shown here), and increase gradually with greater step size. The B-DRUNet also exhibits artefacts, amplifying some structures in the top central part of the image.
@@
Lastly, Figure 6 shows the pixel-wise posterior standard deviation, as calculated with the LMMO, Prox-DRUNet, GS-DRUNet and B-DRUNet denoisers. For reference, next to each posterior standard deviation plot, we report the residual obtained by comparing $\hat{x}_{MMSE}$ to the true image (these standard deviations represent the models’ marginal predictions for these residuals, at the pixel level). We observe that LMMO, Prox-DRUNet, GS-DRUNet produce uncertainty plots that are broadly in agreement with their residuals. For Prox-DRUNet and LMMO, uncertainty concentrates around edges and contours, as expected for a deconvolution problem, whereas the uncertainty estimates of GS-DRUNet are aligned with its reconstruction artefacts. Conversely, B-DRUNet produces uncertainty plots that highlight homogenenous regions, and which do not align well with its residual. We conclude that Prox-DRUNet is the most appropriate denoising architecture for Bayesian PnP inference in Poisson image deblurring problems and use this as the prior for the rest of our experiments.
@@ 4.3.1 Experimental set up
We now compare different PnP Langevin strategies with Prox-DRUNet as prior. The primary purpose of this experiment is to provide an empirical and qualitative exploration of the stability boundaries and performance trade-offs of each algorithm. Note that the considered algorithm often demonstrate stability well beyond the range of step sizes predicted by the theory available (an effect we attribute in part to pessimistic stability estimates, to the algorithm’s constraints, and the use of a mirror map), hence we aim to characterize their behavior as the step size is further increased. Specifically, we study how different choices of step size affects convergence speed, reconstruction quality, and numerical stability. The results are intended to serve as an illustrative guide demonstrating the visual cues of both stable (well-chosen large step sizes) and unstable (excessively large step sizes) regimes, such as the emergence of artifacts or over-smoothing. We compare MMSE reconstructions computed with different sampling algorithms for a range of step sizes, both quantitatively and qualitatively, to derive practical insights for each algorithm. Additionally, we compare the convergence speed between different sampling algorithms for the recommended step size over a set of images by looking at mean cumulative quality metrics. How image quality and convergence speed interrelate for different algorithms can guide the choice of sampling algorithm according to given priorities with regards to metric and computational budget. Finally, we illustrate the mixing properties of these algorithms by looking at the autocorrelation of the corresponding Markov chains.
@@
For this experiment, we use images from the set3c data set, depicted in Figure 3, the blur kernel displayed Figure 2(b) and photon level $\alpha=20$. We choose the starfish image to illustrate qualitative results. With regards to the algorithms and their step sizes, for the RPnP-ULA (Algorithm 1) and PPnP-ULA (Algorithm 2), we test a range of step sizes $\delta=c\cdot\delta_{L}$, with $c\in[1,10^{3}]$. For RPnP-SKROCK (Algorithm 3), we test step sizes of the form $\delta=c\cdot\ell_{s}\delta_{L}$, with $c\in[0.1,10^{2}]$, where $\ell_{s}=(s-0.5)^{2}(2-4/3\eta)-1.5$ with $s=10$ and $\eta=0.05$; for more details see [50]. For the PnP-MLA (Algorithm 4), we test step sizes of the form $\delta=c\cdot\delta_{L}$, where $c\in[10,10^{3}]$. The Prox-DRUNet denoiser parameter $\epsilon$ is set to $\epsilon=(20/255)^{2}$ for RPnP-ULA, PPnP-ULA and RPnP-SKROCK, whereas we use $\epsilon=(25/255)^{2}$ for PnP-MLA (see Appendix B.2 for complementary results for PnP-MLA with $\epsilon=(20/255)^{2}$). For the Euclidean algorithms we set $\beta$ to $1\%$ of the mean intensity value of the observed data. In all cases, to improve numerical stability and reduce PnP artefacts, we enforce equivariance to horizontal and vertical flips as well as to rotations of multiples of 90 degrees by randomization.
(a) butterfly
(b) leaves
(c) starfish
Figure 9: Comparing the auto-correlation function (ACF) for PnP-MLA, RPnP-ULA, PPnP-ULA and RPnP-SKROCK for the set3c data set.
@@ 4.3.2 Experimental results
In Figure 8.3, we observe that RPnP-SKROCK is stable as the step size increases and does not show artefacts in its reconstructions. For large step sizes, however, we see significant oversmoothing which is a sign of excessive bias, see e.g. Figure 8.3(d) and 8.3(e). In this case, PSNR and SSIM continue to improve, while LPIPS worsens, quantitatively confirming a loss of detail. The overall performance deteriorates for a step size for which $c>0.3\cdot 10^{2}$. Therefore, for this specific example, we would recommend to choose a step size for which $c=0.02\cdot 10^{2}$ or $c=0.3\cdot 10^{2}$ depending on whether objective (PSNR, SSIM) or perceptive quality (LPIPS) is more important. More generally, we advise to check both for qualitative cues such as artifacts or excessive smoothing, and quantitative metrics such as convergence speed, to determine an appropriate step size for the problem at hand. Moreover, in Figure 8.4, we observe that PnP-MLA performs relatively similarly for all the considered step sizes, consistently recovering fine detail without any significant oversmoothing effect.
In order to illustrate the mixing properties of the presented algorithms, Fig. 9 shows the autocorrelation function (ACF) of the slowest and fastest mixing components [^10] The slow and fast mixing components were identified by selecting the pixel trajectories with the maximum and minimum empirical variance across the chain, respectively. The ACF was computed using the Fast Fourier Transform (FFT). of the Markov chains for images of the set3c dataset. Efficient sampling algorithms generate samples that are not highly correlated, which can be observed by a rapid decay of the ACF curve towards zero as the lag increases. The analysis of the slowest component highlights the most significant performance differences. The RPnP-SKROCK (lilac) is often most efficient, showing a very rapid decay in correlation. In contrast, the decay of PnP-MLA, PPnP-ULA and RPnP-ULA varies per image, with a decay within 15-30 lags. For the fast-mixing component, all algorithms perform well, with correlations dropping to near-zero almost immediately (within 5 lags). Overall, the RPnP-SKROCK algorithms demonstrate superior performance by effectively handling the more challenging slow-mixing components; the results for the remaining algorithms are mixed but indicate good overall sampling efficiency.
@@
Figure 11: Mean cumulative metrics for the set3c data set for photon level $\alpha=20$ using selected time steps, with x-axis in log scale to show the tail behaviour more clearly.
Finally, in Figure 11 we compare quality metrics (PSNR and LPIPS) against neural function evaluations (NFE) averaged over set3c using recommended step sizes to assess the convergence speed. We observe that RPnP-SKROCK outperforms all other algorithms in terms of PSNR in performance and convergence speed. In addition, RPnP-ULA and PPnP-ULA show similar performance and convergence speed while they outperform RPnP-SKROCK and PnP-MLA in terms of LPIPS. These differences relate to the fact that the two algorithms target two slightly different invariant distributions, in particular PnP-MLA targeting (1) and RPnP-SKROCK targeting (8). In addition, some further bias is introduced by the fact that both algorithms are operating with a finite step-size $\delta$, see for example the discussion in [1]. PnP-MLA shows slower convergence in terms of PSNR, but it performs similarly with the other algorithms in terms of LPIPS with sharp convergence in $\sim 10^{4}$ NFEs.
@@ 4.4.2 Experimental results
4.5 Low-dose computer tomography experiment
4.5.1 Experimental set up
We now present an experiment on low-dose Computed Tomography (CT) reconstruction, a canonical inverse problem commonly encountered in medical imaging and non-destructive testing. In the low-dose setting, CT measurements exhibit strong Poisson noise, making it an ideal test case for the proposed algorithms. We consider a standard parallel-beam CT configuration, where the forward operator $A$ is a discretization of the Radon transform in a full-view setup. While this full-view setup avoids sparse-view and limited-angle artifacts, the problem remains severely ill-conditioned and hence requires significant regularization to produce solutions that are well posed. To simulate low-dose acquisitions, we apply this operator to ground truth lung images from the the LIDC-IDRI dataset [5] and then corrupt the resulting sinograms with aggressive Poisson noise, effectively limiting the photons per view. For the PnP prior, we trained a new GS-DRUNet denoiser $D_{\epsilon}$ specifically for this task. The network was trained on 2D slices from the LIDC-IDRI training set, following the data preparation and training strategy of [68] and [31]. We set the photon level $\alpha=10$, the regularization parameter $\rho=4$, the denoiser level $\epsilon=(20/255)^{2}$, and determined appropriate step sizes for PnP-MLA ($\delta_{MLA}=10^{-4}$) and RPnP-SKROCK ($\delta_{SKR}=5\cdot 10^{-5}$) via a grid search. For RPnP-SKROCK, which requires a Lipschitz-continuous likelihood, we used the regularized Poisson likelihood (8) as in Sections 4.3, with $\beta$ is set to $1\%$ of the mean intensity value of the observed data. In contrast, PnP-MLA uses the same likelihood, however with a much smaller $\beta=10^{-8}$ which is required to ensure numerical stability.
4.5.2 Experimental results
Figure 14 presents an example of a reconstruction result for an image from the LIDC-IDRI test set, alongside the standard Filtered Back-Projection (FBP) for reference. Both RPnP-SKROCK and PnP-MLA successfully recover high-quality images from the low-dose CT measurements, demonstrating a significant improvement over FBP. We also show the joint marginal standard deviations for pixel scale and groups of pixels of size $2\times 2$, $4\times 4$ and $8\times 8$ pixels, which display the uncertainty about structures at different sizes and in different regions of the reconstruction. These maps provide valuable insights for interpretation, highlighting uncertainty in areas rich in fine anatomical structures.
Refer to captionRefer to captionRefer to captionRefer to captionGTFBPRefer to captionRefer to captionRefer to captionRefer to captionRPnP-SKROCK(29.2,0.10)PnP-MLA(29.8,0.15)RPnP-SKROCK $\hat{x}_{MMSE}$Std. dev. $1\times$ 1Std. dev. $2\times$ 2Std. dev. $4\times$ 4Std. dev. $8\times$ 8Refer to captionRefer to captionRefer to captionRefer to captionPnP-MLAStd. dev. $1\times 1$Std. dev. $2\times 2$Std. dev. $4\times 4$Std. dev. $8\times 8$
Figure 14: Results of the low-dose CT experiment for $\alpha=10$. First row: Ground truth, FBP, $\hat{x}_{MMSE}$, with reported metrics (PSNR, LPIPS) for two select methods (RPnP-SKROCK and PnP-MLA). Second and third row: The marginal standard deviations of the pixel-wise scale $1\times 1$, and for groups of pixels of size $2\times 2$, $4\times 4$ and $8\times 8$ pixels, for both methods.
5 Discussion
A natural question arising from this work is whether the explicit acceleration of SKROCK could be combined with the geometric adaptation of Mirror Langevin (MLA) to create a further improved sampler. While this presents an intriguing prospect, we understand that these two methodologies represent fundamentally different, and potentially conflicting, paradigms for addressing ill-conditioned problems. SKROCK provides explicit acceleration by using a sophisticated integrator to improve convergence rates as a function of the Euclidean condition number $\kappa$. In contrast, MLA provides implicit acceleration by reconditioning the problem via a mirror map, using a geometry where the effective condition number $\tilde{\kappa}$ is significantly smaller, thus allowing a simpler integrator to perform efficiently. Combining these approaches is non-trivial and opens significant theoretical questions. It is unclear what additional benefit an explicit $\mathcal{O}(\kappa)$-type acceleration scheme would provide if the mirror map has already driven the effective condition number $\tilde{\kappa}\approx 1$. Furthermore, defining and analyzing Nesterov-style momentum within the dual space geometry of a mirror map, particularly for a stochastic process, is highly non trivial. In particular, this extension would require extending to the SDE setting the theories of accelerated mirror descent from optimization [32] and continuous-time ODE perspectives on non-Euclidean acceleration [16, 3]. Given these complexities, we consider the development of a provably accelerated mirror Langevin algorithm a research direction in its own right, and an interesting avenue for future work. Our contribution rigorously studies the practical performance of these two distinct state-of-the-art approaches for the challenging problem of PnP Poisson imaging.
@@ 6 Conclusion
This paper addressed an important gap in Bayesian imaging methodology by making advanced Plug-and-Play (PnP) MCMC a viable and powerful tool for Poisson inverse problems. We systematically adapted and extended state-of-the-art Langevin samplers to handle the poor regularity and constraints of the Poisson likelihood. We focused on two algorithms: the explicitly accelerated RPnP-SKROCK, using reflections and a likelihood approximation, and PnP-MLA, the first PnP mirror Langevin algorithm for imaging, which leverages a problem-adapted geometry to simultaneously deal with the constraints and with the Poisson likelihood without approximations.
Our main contribution is a comprehensive empirical study comparing these distinct geometric and accelerated MCMC methodologies. Using Poisson image deblurring as a testbed, we evaluated a wide range of denoiser architectures, concluding that Prox-DRUNet provides a robust prior for this problem class. The extensive numerical experiments demonstrated that in challenging low-photon settings (small $\alpha$), the proposed sampling algorithms significantly outperform optimization-based PnP strategies in reconstruction accuracy while also providing uncertainty quantification. The results yielded a set of practical recommendations: for low-photon imaging problems, we advocate for PnP-MLA in applications that require superior perceptual quality (LPIPS) and hence a more faithful reconstruction of fine detail and small structures with edges and contours. Conversely, we expect RPnP-SKROCK to deliver a stronger MMSE performance (PSNR) with a comparatively lower computational cost. We expect both methods to produce similarly detailed uncertainty quantification results.
@@
These methodologies also open several avenues for future research. An important direction is to embed Markov kernels like PnP-MLA and RPnP-SKROCK within empirical Bayesian strategies for automatic parameter calibration in semi-blind problems [73, 45]. In addition, while in this paper we have focused on relatively conventional PnP denoiser architectures, we see great potential in using a PnP prior derived from a denoising diffusion model (DM) within a Langevin sampling scheme, like in [44]. Extending the DM-based PnP-ULA of [44] to Poisson imaging problems by using either PnP-MLA or RPnP-SKROCK is also an important perspective for future work. Lastly, a highly promising but more challenging direction for further research would be to explore accelerated PnP-MLA sampling, combining the strengths of PnP-MLA and RPnP-SKROCK.
Finally, we acknowledge the computational cost of the proposed MCMC methods. Unlike optimization techniques that yield a single point estimate, MCMC samplers inherently require a significant number of iterations to explore the posterior distribution for uncertainty quantification, particularly in challenging low-photon regimes. However, there are several promising directions to mitigate this cost. The development of explicitly accelerated mirror Langevin schemes, as mentioned above, could reduce the required number of iterations. Furthermore, the cost per iteration could be lowered by replacing large denoisers like Prox-DRUNet with more lightweight yet powerful architectures, such as those based on model distillation [71] or learned (weakly) convex regularizers [24, 25]. Scalability could also be significantly improved by adapting divide-and-conquer MCMC strategies [74] to imaging inverse problems. These approaches represent important next steps toward making advanced Bayesian inference both powerful and computationally practical for a broader range of applications.
@@
Figures 15(b) to 15(f) show the posterior mean $\hat{x}_{MMSE}$, as calculated for the different denoisers and different algorithms using $10^{6}$ iterations to stress-test the stability of the algorithms. Prox-DRUNet proved to be the most competitive in terms of stability and performance among all the considered denoisers. More specifically, we observe that PPnP-ULA with the LMMO denoiser produces an estimate of the mean with excessive smoothing, while with the GS-DRUNet exhibits checkerboard-like and stripe-like reconstruction artefacts on large parts of the starfish and the background. The B-DRUNet denoiser progressively develops artefacts after approximately $10^{5}$ iterations which deteriorate gravely over time, see Figure 15(f).